# Visual Planning: Let’s Think Only with Images

### v2

@@ 4 Discussions and Analysis
Error Analysis and Case Study. We conduct error analysis for language-based planning and visual planning. We observe that textual planning systems with both SFT and RL are prone to errors when grounding visual inputs to verbalized descriptions during the inference process, with 25.7% of generated coordinate-based layout descriptions and 22.3% of generated ASCII-based representations being mismatched with ground-truth layouts. Qualitative analysis of response from textual RL baselines (Figure 9 in Appendix C.2.1) and proprietary models (Figure 4) also reveal similar observations. Taken together, these results demonstrate an inherent modality gap where language may not be the most accurate and effective representation for vision-first problem. For visual planning, Figure 3 presents visual planning traces generated by LVM across different tasks. We observe that the model occasionally takes non-optimal actions that deviate from the shortest path (FrozenLake example). Surprisingly, VPRL demonstrates the ability to take detours to bypass the obstacles while still progressing towards the goal, whereas VPFT lacks this flexibility and gets stuck, as shown in Figure 4. Additional traces covering optimal, non-optimal, and invalid cases can be found in Appendix C.6. Beyond these in-domain analyses, we further evaluate generalization on larger unseen grids and perturbed image inputs, with results reported in Appendix C.4.
@@ 5 Related Work
MLLM Reasoning
Recent work has extended CoT prompting [55] to MLLMs through approaches such as grounding visual inputs into symbolic representations, such as graphs or bounding boxes [63, 31]. Other approaches integrate tools to generate visualizations during reasoning [24, 65]. For example, o3 model [40] incorporates visual rationales using tools such as cropping and zooming. MVoT [34] is also essentially a form of tool use: instead of relying on external modules, it invokes itself to generate visualizations of textual reasoning. These methods primarily conduct reasoning in language, with visual components merely illustrating the textual rationale rather than serving as the medium of reasoning. In this work, we take a step further to explore whether multi-step planning can emerge purely within visual representations, enabling reasoning without relying on language at all.
@@
Reinforcement Learning for Visual Reasoning
Reinforcement learning has been applied across a wide range of vision-related tasks, especially given the rise of GRPO as in DeepSeek-R1 [17]. Concurrently, in object detection, visual perception [58] is optimized though rewarding high Intersection-over-Union (IoU) scores between predicted and ground-truth bounding boxes [49]. For visual reasoning tasks such as Visual Question Answering (VQA), GRPO has been utilized to optimize the models for longer, more coherent, and logically grounded reasoning traces in textual responses [37, 64, 62, 50]. More recently, similar methods have also been applied to image generation tasks for recursive refinement with textual instructions [18, 52, 27]. These approaches focus on pixel-level fidelity and semantic alignment with text, whereas our work leverages RL for goal-oriented visual planning, optimizing multi-step decision-making through visual state transitions without any reliance on language.
Action-conditional Generative Models
Action-conditional generative models has focused on constructing latent representations of the world and predicting future observations conditioned on given actions [20, 5]. These models learn transition dynamics and are central to model-based reinforcement learning, where they allow agents to simulate potential outcomes without interacting directly with the environment [21]. While effective for representation learning and short-horizon prediction, action-conditional generative models do not perform planning and must therefore be coupled with an external planner. In contrast, our approach constitutes a holistic planner that internalizes planning within the visual generative flow, which is more effective for visual tasks than traditional text-based planners that suffer from a modality gap. It can also benefit from action-conditional generative models by using predicted observations as inputs.
@@ 6 Conclusion
Appendix A Limitations and future work
@@
In this work, we focus exclusively on Large Vision Model (LVM) to investigate visual planning capabilities by eliminating language as a confounding factor for research purposes. As such, this choice constraints the model size to 7B as the only available size of LVM, and excludes recently released native multimodal models capable of generating multimodal outputs [10, 56]. However, we argue that the visual planning paradigm can be extended to broader multimodal generation models for use in more diverse tasks, combined with more modalities, as long as they support image generation.
@@
Additionally, explicitly generating images introduces computational overhead during inference compared to a textual response. However, we argue that language-based reasoning performs worse than visual planning and can be equally or even more time-consuming, especially for thinking models [14]. In our demonstration, Gemini generated over 7,000 thinking tokens yet failed to provide the correct answer in the end. The computation overhead introduced by image generation can be alleviated through more compact image representations using fewer tokens [11], which we advocate for future research.
@@
Another limitation in this work lies in the implementation of dynamics interpreter. For simplicity, we adopt the rule-based approach that compares pixel-wise features between the current state and the previous state (details in Appendix B.3). While effective in our controlled setup, broader task settings with more diverse types of dynamic interpreters are yet to be explored. Nevertheless, we argue that the core idea is extensible and could be supported by either dynamic models to elicit actions from pairs of images (e.g. segmentation [45] or contour detection [35]) or a holistic neural model (e.g. Gemini [14]) that directly judges whether the transitions are valid without explicitly inferring actions. We encourage future research to explore more robust and scalable designs for the visual dynamics interpreter to advance visual planning systems.
Broader Impact
@@
This work introduces a novel paradigm of visual planning, where agents reason and act entirely within the visual modality without reliance on textual intermediaries. By demonstrating that models can plan through sequences of images, this research opens new possibilities for the way human and AI system interacts, particularly in domains like robotics, navigation, and assistive technologies, where perception and decision-making are tightly coupled. As the first step toward planning grounded purely in visual representations, our work lays the foundation for AI systems that integrate both verbal and non-verbal reasoning. We advocate for future research into more holistic multimodal thinking systems where interleaved text and image traces enable richer, more human-like reasoning, and emphasize the importance of strengthening the visual component in such traces for improved planning and cognition.
Appendix B Implementation details
B.1 Dataset
@@
  • SFT in Text (Baseline): For each environment, we sample an optimal trajectory consisting of a sequence of visual states $(v_{0},\ldots,v_{n})$ as the ground truth. Each transition between states is determined by an action, enabling us to derive a corresponding verbalized action sequence $(a_{0},\ldots,a_{n-1})$. The input to the model is formulated by concatenating a textual prompt with an image representation of the initial state $v_{0}$, while the target output is the verbalized action sequence representing the optimal trajectory. We further ablate different variants of the baseline with various representations and tuning methods (SFT and RL) in Appendix C.2. The detailed prompts for all variants are provided in Appendix D.
@@
  • VPFT*: We conduct an ablation study (indicated with *) where VPFT is also trained in two stages, mirroring the structure of VPRL. Stage 1 follows the same procedure as VPRL Stage 1, focusing on format supervision using enumerated visual inputs. Stage 2 reuses the original VPFT training pipeline, learning from optimal trajectories. Experimental results and analysis see Appendix C.5.
@@
B.2 Models
Large Vision Model (LVM) [4] is an autoregressive models for image generation, which is only pretrained with image sequences with no exposure to language data. The model uses a tokenizer based on the VQGAN architecture [13], which extracts visual information from raw images and encodes it into 256 tokens from a fixed codebook. The image is generated in an auto-regressive manner with discrete tokens, which are then fed into the image detokenizer. Although LVM supports multiple model sizes, only the 7B-parameter version is publicly available; thus, we use this variant in our experiments. For a fair comparison, we use Qwen 2.5-VL-Instruct [3] with a matching parameter size as our language-based baseline.
@@ B.3 Reward Implementation
Our method and reward modeling approach are readily generalizable to other visual tasks. With reference to computer vision techniques such as segmentation [45] and contour detection [35], the pixel-level analysis used in our framework can be easily extended to a wide range of structured visual environments. Furthermore, our reward design is broadly applicable to planning tasks in general. Since actions in most planning settings can naturally be categorized into one of three types (valid and helpful, valid but non-progressing, or invalid), our simple reward structure remains intuitive and effective across tasks.
B.4 Training details
@@
Supervised Fine-Tuning (SFT) in Text. In this baseline, planning is formulated in the language modality. Instead of generating an intermediate visual consequence of an action, the model produces a textual description of the intended action sequence. Formally, given an visual input state $v$ and a textual prompt $p$, which represents the task description, the model is trained to generate a verbalized action sequence $t=(t_{1},\ldots,t_{L})$, where each token $t_{i}\in\mathcal{V}_{\text{text}}$ represents an action. The input to the model is the concatenation of the prompt tokens and the visual tokens, and the target is the corresponding action sequence. Following prior work on supervised fine-tuning (SFT) [53] in autoregressive models, we minimize the cross-entropy loss for action prediction:
@@
For all post-training experiments, we apply Low-Rank Adaptation (LoRA) [23] on both attention layers and feed-forward layers. The detailed hyper-parameters are shown in Table 5. Only the loss of the targets is calculated in an instruction-tuning manner [53] for SFT. The image tokenizer and detokenizer are frozen during training. We use the AdamW optimizer [38] for all training procedures.
@@ C.1 VPRL Training
C.2 Trained textual baselines and reward design
@@
Model / Variant EM (%) PR (%) L3 L4 L5 L6 Avg. L3 L4 L5 L6 Avg. Qwen 2.5-VL-Instruct-3B   xxx- SFT   xxxxxx- Direct 87.2 72.0 48.8 28.0 59.0 89.4 84.3 71.3 60.1 76.3   xxxxxx- w/ Coordinates 87.2 78.0 64.8 30.8 65.2 89.6 82.5 74.0 57.2 75.8   xxxxxx- w/ ASCII 79.6 75.6 58.8 34.0 62.0 83.3 82.5 74.8 59.1 74.9   xxx- GRPO   xxxxxx- w/ VPRL progress reward 69.2 52.8 41.2 26.0 47.3 73.6 72.2 66.2 55.0 66.8   xxxxxx- w/ PR metric reward 70.8 60.0 41.6 23.6 49.0 75.5 76.1 65.7 56.0 68.4 Qwen 2.5-VL-Instruct-7B   xxx- SFT   xxxxxx- Direct 97.6 86.0 56.4 34.4 68.6 98.1 92.1 78.9 68.4 84.4   xxxxxx- w/ Coordinates 93.2 88.0 74.8 41.6 74.4 94.1 89.7 81.5 65.5 82.7   xxxxxx- w/ ASCII 93.2 86.0 68.0 45.2 73.1 94.1 88.6 81.3 69.6 83.4   xxx- GRPO   xxxxxx- w/ VPRL progress reward 72.4 64.0 50.4 30.8 54.4 76.2 76.3 69.2 57.8 69.9   xxxxxx- w/ PR metric reward 82.8 68.8 51.6 37.2 60.1 84.9 79.6 71.5 61.0 74.3 LVM-7B   xxx- VPFT (ours) 92.0 82.8 68.8 58.0 75.4 93.1 84.7 73.4 66.9 79.5   xxx- VPRL (ours) 97.6 95.6 90.8 82.4 91.6 98.4 96.0 93.0 85.6 93.2
@@
The example input-output formats for different text-based reasoning variants are shown in Figure 8 in Appendix C.2.1.
@@
RL-trained text baseline. We also explore the feasibility of applying RL to improve the planning performance with multimodal input, given the success of RL in the pure language planning domain [17]. We train an RL-based text model using Qwen-2.5-VL-Instruct-3B and 7B with the GRPO algorithm, with output format shown in Figure 8.
@@
Input Example Refer to caption Task: Generate the shortest valid sequence of actions that moves the elf from the starting position to the goal without stepping into any ice holes (see detailed prompts in Appendix D).
SFT (w/ Coordinates) After analysis, the layout of the given image is as follows:
- Starting position: (3, 3)
- Goal: (2, 2)
- Ice holes: (1, 1), (2, 3)
- Passable cells: (1, 2), (1, 3), (2, 1), (3, 1), (3, 2)
Based on the above layout, the final answer is: <ANSWER>left up</ANSWER>
SFT (w/ ASCII) The symbols used to represent the grid are:
- S denotes the starting position, G the goal, H the ice hole, F passable cells.
HFH
FGH
FFS
Based on the above layout, the final answer is: <ANSWER>left up</ANSWER>
GRPO <think>
1. The elf is in the bottom right cell.
2. The goal is in the top center cell.
3. The top left cell has an ice hole, so the elf cannot go left there.
4. The elf needs to move one cell left to the center cell.
5. Then the elf needs to move one cell up to the goal cell.
The sequence of actions are:
1. Move one cell left to the center cell.
2. Move one cell up to the goal cell.
</think>
<answer>left up</answer>
@@
Level 3 Refer to caption <think>
1. The elf is at the center cell.
2. The goal is the gift box in the top left corner.
3. The elf needs to move 1 cell left and 1 cell up.
4. There are ice holes directly below and to the right, so they are impassable.
</think>
<answer>left up</answer>
Level 4 Refer to caption <think>
1. The elf is at the top left section of the grid.
2. The goal is the gift box in the bottom right section.
3. The elf needs to move 3 cells down and 1 cell right.
4. There are ice holes directly below and to the right, so the elf cannot move through them.
</think>
<answer>down down down right</answer>
Level 5 Refer to caption <think>
1. The elf is at the bottom left.
2. The goal is the gift box in the top right.
3. The elf needs to move 3 cells right and 2 cells up.
4. There are ice holes directly above and to the right, so the elf cannot move through them.
</think>
<answer>right right right up up</answer>
Level 6 Refer to caption <think>
1. The elf is at the top right.
2. The goal is the gift box at the bottom left.
3. The elf needs to move 3 cells down and 1 cell left.
4. There are ice holes directly below and to the right of the elf, so we cannot use those paths.
</think>
<answer>left down down down</answer>
@@
C.2.1 Examples of trained Textual variants

### v3

@@ 4 Discussions and Analysis
Error Analysis and Case Study. We conduct error analysis for language-based planning and visual planning. We observe that textual planning systems with both SFT and RL are prone to errors when grounding visual inputs to verbalized descriptions during the inference process, with 25.7% of generated coordinate-based layout descriptions and 22.3% of generated ASCII-based representations being mismatched with ground-truth layouts. Qualitative analysis of response from textual RL baselines (Figure 9 in Appendix F.2.1) and proprietary models (Figure 4) also reveal similar observations. Taken together, these results demonstrate an inherent modality gap where language may not be the most accurate and effective representation for vision-first problem. For visual planning, Figure 3 presents visual planning traces generated by LVM across different tasks. We observe that the model occasionally takes non-optimal actions that deviate from the shortest path (FrozenLake example). Surprisingly, VPRL demonstrates the ability to take detours to bypass the obstacles while still progressing towards the goal, whereas VPFT lacks this flexibility and gets stuck, as shown in Figure 4. Additional traces covering optimal, non-optimal, and invalid cases can be found in Appendix F.6. Beyond these in-domain analyses, we further evaluate generalization on larger unseen grids and perturbed image inputs, with results reported in Appendix F.4.
@@ 5 Related Work
MLLM Reasoning. Recent work has extended CoT prompting (Wei et al., 2022c) to MLLMs through approaches such as grounding visual inputs into symbolic representations, such as graphs or bounding boxes (Zhang et al., 2024b; Lei et al., 2024). Other approaches integrate tools to generate visualizations during reasoning (Hu et al., 2024; Zhou et al., 2024). For example, o3 model (OpenAI, 2025) incorporates visual rationales using tools such as cropping and zooming. MVoT (Li et al., 2025) is also essentially a form of tool use: instead of relying on external modules, it invokes itself to generate visualizations of textual reasoning. These methods primarily conduct reasoning in language, with visual components merely illustrating the textual rationale rather than serving as the medium of reasoning. In this work, we take a step further to explore whether multi-step planning can emerge purely within visual representations, enabling reasoning without relying on language at all.
@@
Reinforcement Learning for Visual Reasoning. Reinforcement learning has been applied across a wide range of vision-related tasks, especially given the rise of GRPO as in DeepSeek-R1 (Guo et al., 2025a). Concurrently, in object detection, visual perception (Yu et al., 2025) is optimized though rewarding high Intersection-over-Union (IoU) scores between predicted and ground-truth bounding boxes (Shen et al., 2025). For visual reasoning tasks such as Visual Question Answering (VQA), GRPO has been utilized to optimize the models for longer, more coherent, and logically grounded reasoning traces in textual responses (Liu et al., 2025; Zhou et al., 2025; Zhang et al., 2025b; Team et al., 2025). More recently, similar methods have also been applied to image generation tasks for recursive refinement with textual instructions (Guo et al., 2025b; Wang et al., 2025; Jiang et al., 2025). These approaches focus on pixel-level fidelity and semantic alignment with text, whereas our work leverages RL for goal-oriented visual planning, optimizing multi-step decision-making through visual state transitions without any reliance on language.
Action-conditional Generative Models. Action-conditional generative models has focused on constructing latent representations of the world and predicting future observations conditioned on given actions (Ha & Schmidhuber, 2018; Ball et al., 2025). These models learn transition dynamics and are central to model-based reinforcement learning, where they allow agents to simulate potential outcomes without interacting directly with the environment (Hafner et al., 2019). While effective for representation learning and short-horizon prediction, action-conditional generative models do not perform planning and must therefore be coupled with an external planner. In contrast, our approach constitutes a holistic planner that internalizes planning within the visual generative flow, which is more effective for visual tasks than traditional text-based planners that suffer from a modality gap. It can also benefit from action-conditional generative models by using predicted observations as inputs.
@@ 6 Conclusion
Acknowledgements
The work has been supported by the UK Research and Innovation (UKRI) Frontier Research Grant EP/Y031350/1 (the UK government’s funding guarantee for ERC Advanced Grants) awarded to Anna Korhonen at the University of Cambridge. The work has also been supported in part by a Royal Society University Research Fellowship (no 221137; 2022-) awarded to Ivan Vulić, and by the UK EPSRC grant EP/T02450X/1.
Appendix A The Use of Large Language Models
Large language models (LLMs) were used as general-purpose tools in this work. Specifically, LLMs assisted in polishing the writing to improve clarity and readability.
Appendix B Ethics Statement
Our research adheres to rigorous ethical guidelines. We verified the licenses of all softwares and datasets we used in this study and ensured full compliance with their terms. Furthermore, we have thoroughly assessed the project and do not anticipate any additional potential risks.
Appendix C Reproducibility Statement
Appendix E.1 introduces the datasets in details with statistics and processing procedure. Appendix E.2 introduces models we used in our paper, and Appendix E.3 provides detailed information regarding reward implementation for VPRL method. All hyper-parameters and training details are listed in Appendix E.4 for reproducibility. Appendix E.5 introduces the licences for the data and models we used. Prompting templates are shown in Appendix G. All data and scripts will be released publicly upon acceptance to facilitate reproducibility.
Appendix D Limitations and future work
@@
In this work, we focus exclusively on Large Vision Model (LVM) to investigate visual planning capabilities by eliminating language as a confounding factor for research purposes. As such, this choice constraints the model size to 7B as the only available size of LVM, and excludes recently released native multimodal models capable of generating multimodal outputs (Chern et al., 2024; Wu et al., 2024a). However, we argue that the visual planning paradigm can be extended to broader multimodal generation models for use in more diverse tasks, combined with more modalities, as long as they support image generation.
@@
Additionally, explicitly generating images introduces computational overhead during inference compared to a textual response. However, we argue that language-based reasoning performs worse than visual planning and can be equally or even more time-consuming, especially for thinking models (Gemini, 2025). In our demonstration, Gemini generated over 7,000 thinking tokens yet failed to provide the correct answer in the end. The computation overhead introduced by image generation can be alleviated through more compact image representations using fewer tokens (Choudhury et al., 2024), which we advocate for future research.
@@
Another limitation in this work lies in the implementation of dynamics interpreter. For simplicity, we adopt the rule-based approach that compares pixel-wise features between the current state and the previous state (details in Appendix E.3). While effective in our controlled setup, broader task settings involving more complex visual structures are yet to be explored. Nevertheless, we argue that the underlying reward formulation remains extensible, but the primary challenge lies in defining reliable progress signals as visual transitions become more complex. Such signals could be supported by either dynamic models that elicit actions from pairs of images (e.g. segmentation (Ravi et al., 2024) or contour detection (Linsley* et al., 2020)) or a holistic neural model (e.g. Gemini (Gemini, 2025) or a learned reward model) that directly judges whether the transitions are valid without explicitly inferring actions. Alternatively, trajectory-level rollouts with final success feedback could be leveraged to identify actions that contribute to progress toward successful outcomes, eliminating the requirement for an explicit dynamics interpreter. We encourage future research to explore more robust and scalable designs for interpreting visual transitions to advance visual planning systems.
@@
Broader Impact. This work introduces a novel paradigm of visual planning, where agents reason and act entirely within the visual modality without reliance on textual intermediaries. By demonstrating that models can plan through sequences of images, this research opens new possibilities for the way human and AI system interacts, particularly in domains like robotics, navigation, and assistive technologies, where perception and decision-making are tightly coupled. As the first step toward planning grounded purely in visual representations, our work lays the foundation for AI systems that integrate both verbal and non-verbal reasoning. We advocate for future research into more holistic multimodal thinking systems where interleaved text and image traces enable richer, more human-like reasoning, and emphasize the importance of strengthening the visual component in such traces for improved planning and cognition.
Appendix E Implementation details
E.1 Dataset
@@
  • SFT in Text (Baseline): For each environment, we sample an optimal trajectory consisting of a sequence of visual states $(v_{0},\ldots,v_{n})$ as the ground truth. Each transition between states is determined by an action, enabling us to derive a corresponding verbalized action sequence $(a_{0},\ldots,a_{n-1})$. The input to the model is formulated by concatenating a textual prompt with an image representation of the initial state $v_{0}$, while the target output is the verbalized action sequence representing the optimal trajectory. We further ablate different variants of the baseline with various representations and tuning methods (SFT and RL) in Appendix F.2. The detailed prompts for all variants are provided in Appendix G.
@@
  • VPFT*: We conduct an ablation study (indicated with *) where VPFT is also trained in two stages, mirroring the structure of VPRL. Stage 1 follows the same procedure as VPRL Stage 1, focusing on format supervision using enumerated visual inputs. Stage 2 reuses the original VPFT training pipeline, learning from optimal trajectories. Experimental results and analysis see Appendix F.5.
@@
E.2 Models
Large Vision Model (LVM) (Bai et al., 2024) is an autoregressive models for image generation, which is only pretrained with image sequences with no exposure to language data. The model uses a tokenizer based on the VQGAN architecture (Esser et al., 2021), which extracts visual information from raw images and encodes it into 256 tokens from a fixed codebook. The image is generated in an auto-regressive manner with discrete tokens, which are then fed into the image detokenizer. Although LVM supports multiple model sizes, only the 7B-parameter version is publicly available; thus, we use this variant in our experiments. For a fair comparison, we use Qwen 2.5-VL-Instruct (Bai et al., 2025) with a matching parameter size as our language-based baseline.
@@ E.3 Reward Implementation
Our method and reward modeling approach are readily generalizable to other visual tasks. With reference to computer vision techniques such as segmentation (Ravi et al., 2024) and contour detection (Linsley* et al., 2020), the pixel-level analysis used in our framework can be easily extended to a wide range of structured visual environments. Furthermore, our reward design is broadly applicable to planning tasks in general. Since actions in most planning settings can naturally be categorized into one of three types (valid and helpful, valid but non-progressing, or invalid), our simple reward structure remains intuitive and effective across tasks.
E.4 Training details
@@
Supervised Fine-Tuning (SFT) in Text. In this baseline, planning is formulated in the language modality. Instead of generating an intermediate visual consequence of an action, the model produces a textual description of the intended action sequence. Formally, given an visual input state $v$ and a textual prompt $p$, which represents the task description, the model is trained to generate a verbalized action sequence $t=(t_{1},\ldots,t_{L})$, where each token $t_{i}\in\mathcal{V}_{\text{text}}$ represents an action. The input to the model is the concatenation of the prompt tokens and the visual tokens, and the target is the corresponding action sequence. Following prior work on supervised fine-tuning (SFT) (Wei et al., 2022a) in autoregressive models, we minimize the cross-entropy loss for action prediction:
@@
For all post-training experiments, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2022) on both attention layers and feed-forward layers. The detailed hyper-parameters are shown in Table 5. Only the loss of the targets is calculated in an instruction-tuning manner (Wei et al., 2022a) for SFT. The image tokenizer and detokenizer are frozen during training. We use the AdamW optimizer Loshchilov & Hutter (2019) for all training procedures.
@@ F.1 VPRL Training
F.2 Trained textual baselines and reward design
@@
Model EM (%) PR (%) L3 L4 L5 L6 Avg. L3 L4 L5 L6 Avg. Qwen 2.5-VL-Instruct-3B   xxx- SFT   xxxxxx- Direct 87.2 72.0 48.8 28.0 59.0 89.4 84.3 71.3 60.1 76.3   xxxxxx- w/ Coordinates 87.2 78.0 64.8 30.8 65.2 89.6 82.5 74.0 57.2 75.8   xxxxxx- w/ ASCII 79.6 75.6 58.8 34.0 62.0 83.3 82.5 74.8 59.1 74.9   xxx- GRPO   xxxxxx- w/ VPRL progress reward 69.2 52.8 41.2 26.0 47.3 73.6 72.2 66.2 55.0 66.8   xxxxxx- w/ PR metric reward 70.8 60.0 41.6 23.6 49.0 75.5 76.1 65.7 56.0 68.4 Qwen 2.5-VL-Instruct-7B   xxx- SFT   xxxxxx- Direct 97.6 86.0 56.4 34.4 68.6 98.1 92.1 78.9 68.4 84.4   xxxxxx- w/ Coordinates 93.2 88.0 74.8 41.6 74.4 94.1 89.7 81.5 65.5 82.7   xxxxxx- w/ ASCII 93.2 86.0 68.0 45.2 73.1 94.1 88.6 81.3 69.6 83.4   xxx- GRPO   xxxxxx- w/ VPRL progress reward 72.4 64.0 50.4 30.8 54.4 76.2 76.3 69.2 57.8 69.9   xxxxxx- w/ PR metric reward 82.8 68.8 51.6 37.2 60.1 84.9 79.6 71.5 61.0 74.3 LVM-7B   xxx- VPFT (ours) 92.0 82.8 68.8 58.0 75.4 93.1 84.7 73.4 66.9 79.5   xxx- VPRL (ours) 97.6 95.6 90.8 82.4 91.6 98.4 96.0 93.0 85.6 93.2
@@
The example input-output formats for different text-based reasoning variants are shown in Figure 8 in Appendix F.2.1.
@@
RL-trained text baseline. We also explore the feasibility of applying RL to improve the planning performance with multimodal input, given the success of RL in the pure language planning domain (Guo et al., 2025a). We train an RL-based text model using Qwen-2.5-VL-Instruct-3B and 7B with the GRPO algorithm, with output format shown in Figure 8.
@@
Input Example Refer to caption Task: Generate the shortest valid sequence of actions that moves the elf from the starting position to the goal without stepping into any ice holes (see detailed prompts in Appendix G).
SFT (w/ Coordinates) After analysis, the layout of the given image is as follows: - Starting position: (3, 3) - Goal: (2, 2) - Ice holes: (1, 1), (2, 3) - Passable cells: (1, 2), (1, 3), (2, 1), (3, 1), (3, 2) Based on the above layout, the final answer is: <ANSWER>left up</ANSWER>
SFT (w/ ASCII) The symbols used to represent the grid are: - S denotes the starting position, G the goal, H the ice hole, F passable cells. HFH FGH FFS Based on the above layout, the final answer is: <ANSWER>left up</ANSWER>
GRPO <think> 1. The elf is in the bottom right cell. 2. The goal is in the top center cell. 3. The top left cell has an ice hole, so the elf cannot go left there. 4. The elf needs to move one cell left to the center cell. 5. Then the elf needs to move one cell up to the goal cell. The sequence of actions are: 1. Move one cell left to the center cell. 2. Move one cell up to the goal cell. </think> <answer>left up</answer>
@@
Level 3 Refer to caption <think> 1. The elf is at the center cell. 2. The goal is the gift box in the top left corner. 3. The elf needs to move 1 cell left and 1 cell up. 4. There are ice holes directly below and to the right, so they are impassable. </think> <answer>left up</answer>
Level 4 Refer to caption <think> 1. The elf is at the top left section of the grid. 2. The goal is the gift box in the bottom right section. 3. The elf needs to move 3 cells down and 1 cell right. 4. There are ice holes directly below and to the right, so the elf cannot move through them. </think> <answer>down down down right</answer>
Level 5 Refer to caption <think> 1. The elf is at the bottom left. 2. The goal is the gift box in the top right. 3. The elf needs to move 3 cells right and 2 cells up. 4. There are ice holes directly above and to the right, so the elf cannot move through them. </think> <answer>right right right up up</answer>
Level 6 Refer to caption <think> 1. The elf is at the top right. 2. The goal is the gift box at the bottom left. 3. The elf needs to move 3 cells down and 1 cell left. 4. There are ice holes directly below and to the right of the elf, so we cannot use those paths. </think> <answer>left down down down</answer>
@@
F.2.1 Examples of trained Textual variants